OpenStudy (anonymous):

Find S (long integral sign) (x^1/2 - 4) (x^-1/2 - 1)dx. Show Workout Step by Step Please Thank you

hartnn (hartnn):

can you foil ? (a+b)(c+d) = ... ?

OpenStudy (tkhunny):

Entirely algebraic. What's stopping you from multiplying the two factors and having only exponents you can deal with?

OpenStudy (anonymous):

I can't do the x^1/2 times by the x^-1/2.....what would I get?

hartnn (hartnn):

yeah, after you multiply u get three terms, and you use the fact that integral x^n dx = x^{n+1}/(n+1) +c

hartnn (hartnn):

the \(\huge {x}^{1/2}\times x^{-1/2}=x^{1/2+(-1/2)}=...? \)

OpenStudy (anonymous):

Yeah lol

OpenStudy (tkhunny):

For \(a\ne0\), we have \(a\cdot\dfrac{1}{a} = a^{1}\cdot a^{-1} = a^{1-1} = a^{0} = 1\)

OpenStudy (anonymous):

Would that give me x^0 which would make x^1 or just 1?

hartnn (hartnn):

x^0 =1 yes

OpenStudy (anonymous):

I got it thanks

hartnn (hartnn):

cool :) welcome ^_^

hartnn (hartnn):

and WELCOME to OpenStudy :D

OpenStudy (tkhunny):

You DO have to stay away from x = 0.

OpenStudy (anonymous):

Thanks... Got it @tkhunny and thanks @hartnn. I'll be on here a lot more now haha :D

hartnn (hartnn):

Glad to hear that :)